$SPAD/src/input d01gbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ndim:=4
e:=(4.0*X[1]*X[3]*X[3]*exp(2.0*X[1]*X[3])/((1.0+X[2]+X[4])**2))
f:ASP4(FUNCTN):=retract e
a:Matrix SF:=[[0.0,0.0,0.0,0.0]]
b:Matrix SF:=[[1.0,1.0,1.0,1.0]]
maxcls:=20000
eps:=0.01
lenwrk:=500
wrk:Matrix SF:=new(1,lenwrk,0.0);
mincls:=1000
result:=d01gbf(ndim,a,b,maxcls,eps,lenwrk,mincls,wrk,-1,f)

———-

References^{[1] nothing}

4